OpenStudy (anonymous):

arithmetic where : t10=8 and t18=-40. find the first 3 terms of the sequence.

OpenStudy (anonymous):

8 +8r = -40 r = -6 a +9r = 8 a = 8 +54 = 62 second: 56 third: 50

OpenStudy (anonymous):

could you explain please?

OpenStudy (anonymous):

Well, an arithmetic sequence is one that has a start value and adds a number to that value, so it has the pattern: a, a+r, a+2r, etc... therefore, to find the rate when given two values in the pattern, just use that rule :D